You drop a ball outside of your window, releasing it from rest. After falling 2.50 m, the ball passes the top of a 2.00-m high window on a lower floor. How long does it take to cross the window?


Correction on the give numbers, sorry. After falling 3.50 m, the book passes the top of a 3.00-m high window...
Answer #1

final velocity (at the end of window)

v^2 = 2 g h = 2 *9.8* ( 3.5 + 3)

v = 11.2587 m/s

intial velocity (at the top of window)

u^2 = 2 *9.8* 3.5

u = 8.28 m/s

using 1st equation of motion

v = u + gt

t = (11.2587 - 8.28) / 9.8

t = 0.3 s

=======

Comment in case any doubt, will reply for sure.. Goodluck